Suppose that you have a supply of a 20% solution of alcohol and a 80% solution of alcohol. How many quarts of each should be mixed to produce 90 quarts that is 50% alcohol?
Answer by edjones(8007)

Let x=quarts of 20% solution, 90-x = quarts of 80% solution
.2 * x = 20x
.8 * (90-x) = 72-.8x
.5 * 90 = 45
.
.2x-.8x+72=45
-.6x=-27
x=45 quarts
90-x=45 quarts
.
Ed